OpenStudy (anonymous):

Factor the expression: x^2 - x - 6

OpenStudy (anonymous):

x2 - x - 6? x^2 +2x - 3x -6=0 x(x+2) - 3(x+2)=0 (x+2) (x-3) =0 x+2=0 x-3=0 x=-2 x=3 ={-2, 3} ans

OpenStudy (anonymous):

Thanks before I got it backwards. I got 2 and -3.

OpenStudy (anonymous):

Hahah yeah it happends